# The Hitchhiker’s Guide to Program Analysis, Part II: Deep Thoughts by LLMs

### v2

@@ 5.3. RQ1: Effectiveness
5.3.4. Analysis of FNs.
Despite the number of FN shown in the table 2 is 0, BugLens still produced several FNs but gets mitigated by majority voting. Specifically, we observed that the FNs were concentrated in the Constraint Assessor (ConA) component, and the SecIA component typically does not generate FNs due to the arbitrary control hypothesis (AC-Hypo), which will be discussed in §5.4.2 (we also provide a case study in §A).
@@
    Overlooked Bypass Conditions: The LLM sometimes failed to recognize complex conditions allowing mitigations or checks within the code path to be bypassed, as discussed in §3.3.5.
@@
    Misinterpreting Validation vs. Sanitization: LLMs occasionally misclassified sanitization as validations, thereby missing vulnerabilities, as detailed in §3.4.
@@
Nevertheless, SAG represents guidance, and its effectiveness is not absolute. The degree to which an LLM can accurately follow these complex instructions can vary, especially across different models. Our results (RQ2/RQ3, Table 3) indicate that while SAG provides significant benefits, some models may still struggle to fully leverage the guidance, potentially resulting in some residual FNs.
Takeaway. BugLens can effectively post-refine the results of exisiting static analyzers, hugely improving the precision, and can even find missed bugs.
@@
Model Full Design w/o SAG Baseline
@@ 5.4. RQ2: Component Contribution
    Baseline: This configuration employs the simple prompt design (§3.1.2) without the SecIA or ConA components. It establishes the baseline performance relying primarily on the LLM’s inherent capabilities with minimal guidance. The baseline design provides a starting point for comparison.
@@ 5.4.3. Constraint Assessor (ConA) without Structured Analysis Guidance (SAG)
While this multi-phase workflow (i.e., SecIA + ConA) significantly reduces the high volume of FPs seen in the Baseline; for instance, Claude 3.7 Sonnet’s FPs dropped from 51 to 2, and Gemini 1.5 Pro’s from 24 to 4. It also leads to a significant increase in False Negatives (FNs) for Gemini 2.5 Pro (FN=14), Claude 3.7 Sonnet and DeepSeek R1 (FN=17), and OpenAI o1 (FN=10) in the ‘w/o SAG’. This supports our hypothesis (§3.1.2) that providing constraints, while helpful for pruning obvious non-bugs, can encourage LLMs to become overly confident. Once patterns suggesting data validity are identified, the LLM may default to classifying the issue as “not a bug,” reflecting a potential statistical bias towards common safe patterns rather than performing nuanced reasoning about subtle flaws or bypass conditions.
@@ 5.4.4. Structured Analysis Guidance (SAG)
Takeaway. The design components of BugLens enables effective LLM bug analysis by significantly reducing both FP and FN compared to baseline prompting.
@@ 5.5. RQ3: Model Versatility
Conversely, while the ‘Full Design’ significantly improved the F1 scores for Gemini 2.5 Pro (0.57) and Claude 3.7 Sonnet (0.54) compared to their baseline or ‘w/o sag’ results by drastically cutting down FPs, they still struggled with relatively high False Negatives (FN=12 and FN=13, respectively). This suggests that even with the SAG, these powerful models may face challenges in precisely balancing the various analytical steps or interpreting the nuanced instructions within our workflow, possibly still exhibiting a degree of the previously mentioned over-confidence (for “sanity check”) that SAG could not fully overcome in their case. DeepSeek R1 (F1=0.77) showed a strong, balanced improvement, landing between the OpenAI models and the Gemini/Claude in terms of final performance with the full design. This demonstrates the potential of the BugLens using in open-sourced models.
@@
Takeaway. BugLens shows broad applicability and yields promising results across diverse LLMs, including the open-source DeepSeek R1. OpenAI’s o3-mini currently gets the best result.
@@ 6. Limitation & Discussion
Data Bias. The potential data bias introduced by Suture and CodeQL-SOD could threaten the internal validity. Despite the design of BugLens is completely decoupled from them, we only make tests on top of them might introduce bias into our experiments. BugLens may perform differently on other static analysis tools.
@@
Intergation with More Static Analysis Tools. The CodeQL-SOD is an interesting static analysis tool that is based on the lightweight static analysis framework, CodeQL. Considering that BugLens can effectively improve the precision of existing static analysis tools, we could implement more corsare-grained (with low precision) static analysis tools and integrate them into BugLens. Compared to heavy static analysis tools such as suture, an imprecise static analysis could be much easier to implement and maintain.
@@ 7. Related work
LLMs for bug detection. LLMs have shown promise in bug detection tasks, leveraging their ability to understand code semantics and context. Recent approaches like LLM4SA (wen_automatically_2024, ), LLift (llift, ), and LLMSAN (DBLP:conf/emnlp/WangZSX024, ) combine static analysis with LLMs to refine results or enable end-to-end bug detection. BugLens takes a different approach by leveraging LLMs to reason about complex path constraints and semantic conditions, addressing scenarios where traditional symbolic methods struggle.

### v3

@@ 6.3.3. Analysis of FPs
6.3.4. Analysis of FNs
Despite the number of FN shown in the table 1 is 0, BugLens still produced several FNs but gets mitigated by majority voting. Specifically, we observed that the FNs were concentrated in the Constraint Assessor (ConA) component, and the SecIA component typically does not generate FNs due to the arbitrary control hypothesis (AC-Hypo), which will be discussed in §6.4.2.
@@ 6.3.4. Analysis of FNs
    Overlooked Complex Conditions: The LLM sometimes failed to recognize complex conditions of the program and therefore summarize the pre-/postcondition incorrectly.
@@
    Misinterpreting Validation vs. Sanitization: LLMs occasionally misclassified sanitization as validations, thereby missing vulnerabilities.
@@
Nevertheless, the impact of SAG is not absolute. The ability of a language model to follow such complex instructions can vary significantly, particularly across different models. Our results in RQ2/RQ3 (Table 2) indicate that some models could still cause some residual FNs even with the SAG.
Takeaway 1. BugLens can effectively post-refine the results of exisiting static analyzers, hugely improving the precision, and can even find missed bugs.
@@
Model Full Design w/o SAG Simple Prompt
@@ 6.4. RQ2: Component Contribution
    Baseline: This configuration employs the simple prompt design, based on related works (wen_automatically_2024, ; li_iris_2025, ). It establishes the baseline performance relying primarily on the LLM’s inherent capabilities with minimal guidance. The baseline design provides a starting point for comparison.
@@ 6.4.3. Constraint Assessor (ConA) without Structured Analysis Guidance (SAG)
While this multi-phase workflow (i.e., SecIA + ConA) significantly reduces the high volume of FPs seen in the Baseline; for instance, Claude 3.7 Sonnet’s FPs dropped from 51 to 2, and Gemini 1.5 Pro’s from 24 to 4. It also leads to a significant increase in False Negatives (FNs) for Gemini 2.5 Pro (FN=14), Claude 3.7 Sonnet and DeepSeek R1 (FN=17), and OpenAI o1 (FN=10) in the ‘w/o SAG’. This supports our hypothesis (§LABEL:subsubsec:naive-llm) that providing constraints, while helpful for pruning obvious non-bugs, can encourage LLMs to become overly confident. Once patterns suggesting data validity are identified, the LLM may default to classifying the issue as “not a bug,” reflecting a potential statistical bias towards common safe patterns rather than performing nuanced reasoning about subtle flaws or bypass conditions.
@@ 6.4.4. Structured Analysis Guidance (SAG)
Takeaway 2. The design components of BugLens enables effective LLM bug analysis by significantly reducing both FP and FN compared to baseline prompting.
@@ 6.5. RQ3: Model Versatility
Conversely, while the ‘Full Design’ significantly improved the F1 scores for Gemini 2.5 Pro (0.57) and Claude 3.7 Sonnet (0.54) compared to their baseline or ‘w/o sag’ results by drastically cutting down FPs, they still struggled with relatively high false negatives (FN=12 and FN=13, respectively). This suggests that even with the SAG, these powerful models may face challenges in precisely balancing the various analytical steps or interpreting the nuanced instructions within our workflow, possibly still exhibiting a degree of the over-confidence (for “sanity check”) that SAG could not fully overcome in their case. DeepSeek R1 (F1=0.77) showed a strong, balanced improvement, landing between the GPT models and the Gemini/Claude in terms of final performance with the full design. This demonstrates the generality of the BugLens to open-sourced models.
@@
Takeaway 3. BugLens shows broad applicability and yields promising results across diverse LLMs, including the open-source DeepSeek R1. OpenAI’s o3-mini currently gets the best result.
@@ 7. Limitation & Discussion
Dependency on LLMs. The performance of BugLens is influenced by the capabilities of the underlying LLMs used for reasoning. While this dependency could potentially impact external validity, our experiments demonstrate BugLens’s robustness across different models. We have evaluated BugLens with multiple LLMs, including both closed-source and open-source models like Deepseek R1, which achieves approximately 80% of the performance of top-performing models. These results confirm that BugLens’s approach generalizes well across different models, though performance variations exist.
Limited Checkers and Bug Scope. Our evaluation uses only taint-style checkers—those inherited from and Suture (and Dr.checker), plus one CodeQL port—focused on memory-safety and DoS bugs in the Linux kernel. This narrow setup threatens external validity, and the performance we report may not carry over to other bugs (e.g., data races) or analysis frameworks. Although the extra CodeQL checker suggests BugLens can transfer across tooling, a broader study is needed for general applicability.
@@
Intergation with More Static Analysis Tools. The CodeQL-OOB is an interesting static analysis tool that is based on the lightweight static analysis framework, CodeQL. Considering that BugLens can effectively improve the precision of existing static analysis tools, we could implement more corsare-grained (with low precision) static analysis tools and integrate them into BugLens. Compared to heavy static analysis tools such as suture, an imprecise static analysis could be much easier to implement and maintain.
@@ 8. Related work
LLM for Program Analyses & Bug Detection LLMs have been widely applied to program analysis tasks for bug detection. IRIS (li_iris_2025, ) combines LLMs’ contextual understanding with CodeQL queries to enhance taint analysis. LLM4SA (wen_automatically_2024, ) leverages LLMs to refine static analysis results. Their approach serves as a baseline for BugLens (referred as simple prompt). LLMDFA and LLMSAN (DBLP:conf/emnlp/WangZSX024, ; DBLP:conf/nips/WangZSXX024, ) use LLMs to perform data flow analysis and bug detection. LLift (llift, ) focuses on use-before-intialization bugs and prompts LLMs to identify and summarize possible initializers. Focusing on real-world problems, BugLens upgrades the scope to a general taint-style bug detection cross multiple functions.